Please circle the correct anwser
Help with 3,5,6,7,8 thank you
The Earth oelits the san oace evry 365 25 Find the wverage angular speed of Earth Awr in units of 06 100 points okofthe speed ofa.m..E 1.ให e.e bedete. the speed of the belkt 000 (part 2 of2) 10.0 points with constand angular acceleratie What is the speed od the balet? 007 10.0 polnts lar speed o 0.30 re/s in 39.9 what in its anpular acceleration in ral 008 10.0 po The tape s a videotape cwwette has a total tape starts to play, the hall reel has n our AtREr pek¡ด «kring tle play, b.htmb will tbe have the same anglar sped What is this conon angular speed
Answer #1

6

w = angular velocity = 542 rev/min = 542 (2\pi/60) rad/s = 56.73 rad/s

\theta = angular displacement = 41 deg = 41 x 0.0175 rad = 0.7175 rad

t = time for the displacement = ?

Using the equation

\theta = w t

0.7175 = 56.73 t

t = 0.01265 sec

d = distance traveled by the bullet = 99 cm = 0.99 m

v = speed of the bullet = ?

speed is given as

v = d/t

v = 0.99/0.01265

v = 78.3 m/s